OpenStudy (anonymous):

Will fan Simplify 8+12-6+13

OpenStudy (anonymous):

Really?..

OpenStudy (kendricklamar2014):

Hey, this is a very hard problem :P

OpenStudy (studygurl14):

No order of operations here. Just add and subtract right to left

OpenStudy (studygurl14):

left to right I meant, lol

OpenStudy (anonymous):

i only know how to simplfy fractions

OpenStudy (kendricklamar2014):

\[8+12−6+13\] \[=20−6+13\] \[=14+13\] \[= 27\]

OpenStudy (kendricklamar2014):

That's how u solve this problem ^

OpenStudy (ihelpyoulearn):

27! All you have to do is add and subtract :-/